OpenStudy (anonymous):

WILL FAN!! Use the slope formula to find the slope of the line passing through the given points. (2, 1) and (4, –5)

OpenStudy (anonymous):

\(\Huge\sf\color{#00ffff}{Ugh}\) \(\Huge\sf\color{#ff1ff1}{kinda}\) \(\Huge\sf\color{#ff007f}{hard}\)

OpenStudy (anonymous):

\(\Huge\sf\color{#FF033E}{Sorry!:(}\)

OpenStudy (anonymous):

???? then why did u respond

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

do you gt a grid or something

OpenStudy (anonymous):

\[m =\left[\begin{matrix}y2 - y1 \\ x2 -x1\end{matrix}\right]\]

OpenStudy (anonymous):

y2-y1 x2-x1 y=-6 x=2 slope is -3

OpenStudy (anonymous):

thats your answer

OpenStudy (anonymous):

fjsjksdjkdfhdhjdsjalfsgdfh